1. Two balls are drawn in succession without replacement from an urn containing

2 green balls and 6 yellow balls. Let M be the random variable representing the

number of yellow balls.

a. What are the possible values of this random variable (range space)?

b. Construct a probability distribution.


1
Expert's answer
2021-05-17T08:11:52-0400

Given,

2 Green balls and 6 yellow balls

Let G denotes green ball and Y denotes yellow ball

and M be the random variable representing the number of yellow balls


So,

Possible values:




Case1:

when M=2 means the probability of getting 2 yellow balls when we draw out the balls one at a time is:

"P(M_1)=\dfrac{6}{8}\times \dfrac{5}{7}=\dfrac{30}{56}"


Case2:

When M=1 and first ball drawn is yellow and second ball drawn is green

Probability "P(M_2)=\dfrac{6}{8}\times\dfrac{2}{7}=\dfrac{12}{56}"


Case3:

When M=1 and first ball drawn is green and second ball drawn is yellow

Probability "P(M_3)=\dfrac{2}{8}\times\dfrac{6}{7}=\dfrac{12}{56}"


Case4:

When M=0 and both the balls drawn are green

Probability P"P(M_4)=\dfrac{2}{8}\times\dfrac{1}{7}=\dfrac{2}{56}"


Probability distribution: